OpenStudy (anonymous):

How do you find the point on the terminal side of θ =pi/4 that has an x coordinate of 2?

OpenStudy (savvy):

can u plz explain a bit.....i can't get ur question...

OpenStudy (anonymous):

I don't understand it either, this question was given to me on an assignment about trig functions of any angle

OpenStudy (savvy):

ok i think i get it..... i'll explain... gimme a sec

OpenStudy (savvy):

 we need to find the cordinates of the point got it..??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

this is an equilateral triangle, so the \(y\) value is the same as the \(x\) value

OpenStudy (anonymous):

not equilateral, "isosceles"

OpenStudy (anonymous):

angles are \(\frac{\pi}{4}, \frac{\pi}{2},\frac{\pi}{4}\) so two legs are equal

OpenStudy (anonymous):

Im sorry I still do not understand..can you please explain it a little more?